OpenStudy (anonymous):

What value of k makes the statement true? –(k + 12) = 6 + (–12) A. –18 B. –6 C. 6 D. 18

OpenStudy (anonymous):

@SydtheKid913

sammixboo (sammixboo):

Well, you can start this problem by adding \(\rm \color{#20bd23}{6~+~(-12)}\) So what is \(\rm \color{#20bd23}{6~+~(-12)}\)

OpenStudy (anonymous):

umm -6

sammixboo (sammixboo):

Correct, so we have \(\rm \color{#20bd23}{-(k~+~12)~=~-6}\) We can simplify the problem to make the problem look like this {-k ~- ~12~ =~-6}\) Now we add 12 to both sides of the equation :) \(\rm \color{#20bd23}{-k ~- ~12~+~12 =~-6~+~12}\) \(\rm \color{#20bd23}{-k~=~-6~+~12}\) What is -6 + 12?

OpenStudy (anonymous):

6

sammixboo (sammixboo):

Yes, so now we have \(\rm \color{#20bd23}{-k~=~6}\) Now one last step is to divide both sides of the equations by -1 \(\rm \color{#20bd23}{-k\div-1~=~6\div -1}\) \(\rm \color{#20bd23}{k~=~6\div -1}\)

sammixboo (sammixboo):

So what is 6/-1?

OpenStudy (anonymous):

5!

sammixboo (sammixboo):

Mmm no. Try again :P

sammixboo (sammixboo):

6/-1

OpenStudy (anonymous):

i dont know

sammixboo (sammixboo):

A positive number divided by a negative number will always be negative, so what is 6/1? After you have solved that, make the answer negative

OpenStudy (anonymous):

-5?

sammixboo (sammixboo):

Nope.. a number divided by 1 it itself, so 5/1 = 5 3/1 = 3...

OpenStudy (anonymous):

so -3?

sammixboo (sammixboo):

No.. 6/1 = 6, so make that answer negative

OpenStudy (anonymous):

-6?